2 angles are supplementary if their sum is 180 degres. The larger angle measures eight degree more than three times the measure of a smaller angle. If x represents the measure of the smaller angle and these two angles are supplementary, find the measure of each angle. I have the answer but do not know ho to get it 43 degrees & 137 degrees Thank you
Answer by mananth(16949)

2 angles are supplementary if their sum is 180 degres. The larger angle measures eight degree more than three times the measure of a smaller angle. If x represents the measure of the smaller angle and these two angles are supplementary, find the measure of each angle. I have the answer but do not know ho to get it 43 degrees & 137 degrees Thank you


Let the smaller angle be x
The larger angle will be 3x+8
X+3x+8=180
4x=180-8=172
X=172/4
X=43
3*43+8= 137

mananth@hotmail.com